OpenStudy (anonymous):

how to solve this? Lim x + tanx/sinx x -0

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

split it into two terms

ganeshie8 (ganeshie8):

you will need to use the known limit : \(\lim \limits_{x\to 0} \dfrac{\sin x}{x} = 1\)

OpenStudy (anonymous):

is the answer 1/2

OpenStudy (anonymous):

or 2?

ganeshie8 (ganeshie8):

\[\large \lim \limits_{x\to 0} \dfrac{x + \tan x}{\sin x}\]

OpenStudy (anonymous):

Yes it's 2, but do you know how to get it?

OpenStudy (anonymous):

yes i know now

OpenStudy (anonymous):

Alright, cool :)

OpenStudy (anonymous):

thanks anyway